Senior Life Co. is an HMO for businesses in the Portland area. The following account balances appear on the balance sheet of Senior Life Co.: Common stock (800,000 shares authorized; 460,000 shares issued), $5 par, $2,300,000; Paid-in capital in excess of par—common stock, $850,000; and Retained earnings, $34,950,000. The board of directors declared a 2% stock dividend when the market price of the stock was $15 a share. Senior Life Co. reported no income or loss for the current year. Required: A. Journalize the entries to record (1) the declaration of the dividend on December 1, capitalizing an amount equal to market value, and (2) the issuance of the stock certificates on December 31. Refer to the Chart of Accounts for exact wording of account titles. B. Determine the following amounts before the stock dividend was declared: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders’ equity. C. Determine the following amounts after the stock dividend was declared and closing entries were recorded at the end of the year: (1) total paid-in capital, (2) total retained earnings, and (3) total stockholders’ equity.
Answer #1
a1
Stock Dividends 138000 =460000*2%*15
    Stock Dividends Distributable 46000 =460000*2%*5
    Paid in Capital in excess of par-Common Stock 92000
a2
Stock Dividends Distributable 46000
        Common Stock 46000
b
Total paid-in capital 3150000 =2300000+850000
Total retained earnings 34950000
Total stockholders’ equity 38100000
c
Total paid-in capital 3288000 =3150000+138000
Total retained earnings 34812000 =34950000-138000
Total stockholders’ equity 38100000